OpenStudy (anonymous):

Hi there, if anyone could explain this for me, I would be much appreciated, thanks! The world population in the year 2000 was approximately six billion and it will reach seven billion this year (this month probably). Assuming that the population grows at a constant rate (according with p' = rp) calculate the projected world population in 2025. Thanks again, :-)

OpenStudy (anonymous):

increases by 1 billion in 11 years p = 6 + (1/11) t t in years 6 in billions n 2025 p = 6 + (1/11) * 25 = 6 + 2.273 = 8.273 billion approximately

OpenStudy (anonymous):

Thats great, thanks a bunch!!! :-D